Chapter4, Question 2- A commercial jet flies at 550 mi/h for 3000 mi. For how much time does the jet fly? Question 4, A glider flies south east (at 320.0 degrees) at 25.0 Km/h. a wind blows at 12.0 Km/h from 15.0 degrees south of west. What is the new velocity of the glider with respect to the ground in standard position? Question 6- A race car goes from rest to 150 Km/h with an acceleration of 6.0 m/s squared. How many seconds does it take?. Question8- A skateboard starts from rest and accelerates at a rate of 1.30 m/s squared for 3.00s What is his final velocity? Question 10 A train has a final velocity of 110 Km/h . It accelerated for 36 s at 0.50 m/s squared. What was its initial velocity? Question 12-A truck accelerates from rest to 120Km/h in 13 s. Find its acceleration. Question 16- A rock is thrown down with an initial speed of 10.0 m/s from a bridge to the water below. It takes 2.75 s for the rock to hit the water. (a) Find the speed of the rock as it hits the water. (b) How high is the bridge above the water?
Answer #1

Q2: As jet flies at 550mi/hr for 3000 mi, so time of fly = distance/speed = 3000/550 = 5.45 hour Ans

Q6: A the car goes from 0 to 150 km/hr with an acceleration of 6m/s2

Now 150 km/hr = 150X1000/3600 m/s = 41.66 m/s

So from the equation v= u + at, u =0, v=41.66 , a = 6

t = v/a = 41.66/6 = 6.94 seconds Ans

Q8 final velocity v = u + at = 0 + 1.3X3 = 3.9m/s Ans

Q10: v = u + at,

v = 110 km/ hr = 30.5 m/s , t =36s , a = .5

u = v-at = 30.5 - 18 = 12.5 m/s Ans

Q 12 can be solved likewise.

Q16. speed of rock as it hits the water v = u + at = 10 + 2.75X10 = 37.5 m/s

From v2 = u2 + 2gh, h = v2 - u2/2g = 65 meter

Q4:

So the angle between two velocity vector is 180+40+15 = 235

So resultant velocity = (25^2 + 12^2 + 2X25X12Xcos 235)^1/2 = 20.5 km/hr: 40 40 12km/hr 25 km/hr